# PromptHSI: Universal Hyperspectral Image Restoration Framework for Composite Degradation

### v1

@@ Abstract
Recent developments in All-in-One (AiO) RGB image restoration and prompt learning have enabled the representation of distinct degradations through prompts, allowing degraded images to be effectively addressed by a single restoration model. However, this paradigm faces significant challenges when transferring to hyperspectral image (HSI) restoration tasks due to: 1) the domain gap between RGB and HSI features and difference on their structures, 2) information loss in visual prompts under severe composite degradations, and 3) difficulties in capturing HSI-specific degradation representations through text prompts. To address these challenges, we propose PromptHSI, the first universal AiO HSI restoration framework. By leveraging the frequency-aware feature modulation based on characteristics of HSI degradations, we decompose text prompts into intensity and bias controllers to effectively guide the restoration process while avoiding domain gaps. Our unified architecture excels at both fine-grained recovery and global information restoration tasks. Experimental results demonstrate superior performance under various degradation combinations, indicating great potential for practical remote sensing applications. The source code will be publicly released.
Corresponding author, email: cchsu@gs.ncku.edu.tw
@@ 1 Introduction
Figure 1: The All-in-One and universal HSI restoration framework with text-prompt guidance, named PromptHSI, for composite degradation Specifically, informative textual features extracted by pre-trained vision-language model are fed into adapter and then decomposed into an intensity and bias embeddings for degradation-specific restoration with effective cascade guidance.
@@
Hyperspectral images encompass a greater number of spectral bands than RGB images, offering extensive spectral details [44]. Because materials reflect different spectral bands, each substance can have a unique signature for material identification [1, 37, 44, 32]. HSI can capture these valuable information and has broad application value and potential. However, because of the limitation of the device and bandwidth, observed HSIs are usually damaged by different types of degradation from camera sensing and transmission. Thus, HSI restoration is a crucial pre-processing step to recover degraded pixels from flawed observations.
Previous works tended to be based on HSI structures, low-rank sparsity prior [43, 11, 42] for specific degradations such as Gaussian noise [25, 49], stripe effect [24, 22], spatial blurring [33, 19], and cloud occlusion [35, 48]. However, in real-world scenario, HSIs are easily prone to multiple types of degradation, causing these methods to be limited. When HSIs are subject to multiple degradation, some methods based on deep unfolding network [4, 34] or optimization [29, 28, 52, 42] perform significantly below expectations and lack flexibility to use for real-world application.
With the development of AiO RGB image restoration methods [39, 5, 31, 10, 13], which can address multiple degradations in a single network, there is a potential to resolve this situation. However, due to the significant domain gap between the HSI and RGB images, these AiO methods cannot be transferred to HSI restoration. In particular, the visual-prompt approaches [39, 5] are influenced by severe and diverse corruptions. Furthermore, while text-prompts [10, 13] have emerged as an effective approach for controlling All-in-One restoration models, the textual feature representations encoded by widely-used vision-language models (VLMs) pre-trained on RGB-text datasets, such as CLIP [40], exhibit significant domain gaps with HSI features. Given these limitations, leveraging prompt learning frameworks for HSI processing remains an open challenge.
Considered as a whole, the essential questions are Can we leverage prompt learning and vision-language model (VLM) frameworks to address critical challenges in hyperspectral image processing for practical applications, such as achieving composite degradation restoration or controllable restoration through a single unified model? The response is affirmative.
To meet these challenges, we propose PromptHSI, a frequency-aware and prompt-guided framework tailored for HSI’s unique spectral structure and degradation types. Unlike AiO RGB image restoration methods, it ensures spectral fidelity and adaptive degradation response through novel feature modulation strategies. Our universal restoration architecture effectively handles both fine-grained recovery (e.g. spatial-spectral super-resolution) and global information restoration (e.g. decloud, inpainting) within a single model; a frequency-aware feature modulation strategy that decomposes text-prompts into intensity and bias embeddings based on frequency-domain analysis, dispelling the domain gap between VLM and HSI features. In summary, our main contributions can be listed as follows:
  • We present the first universal HSI restoration framework capable of handling composite degradations. Unlike existing HSI restoration methods that suffer from performance drops in composite degradations, our approach maintains consistent high performance across various degradation combinations.
  • We propose a novel text-to-feature modulation mechanism that effectively leverages frequency-domain characteristics of HSI degradations. By decomposing textual features into intensity and bias controller embeddings, our method achieves superior control over the restoration process, outperforming state-of-the-art AiO RGB image restoration methods on composite degradation.
  • Extensive experiments demonstrate that our method not only excels at HSI restoration under various degradation combinations but also exhibits strong generalization to unseen degradation types through flexible text-prompt control. The framework’s effectiveness is thoroughly validated through comprehensive ablation studies and comparisons with existing approaches.
@@
Figure 2: Illustration of the degraded HSI and corresponding description synthesis pipeline. Each degradation is activated with probability $p$. Note that we set $p$ = $0.5$ for all experiments in this paper. To ensure diversity and realistic in the synthetic dataset, each activated degradation selects and applies one of its sub-degradations to the inputted HSI. (More detailed information and setting can be found in Supplementary.)
@@ 2 Related Works
Figure 3: Left, displayed from top to bottom, are HSIs with specific degradation alongside the Fourier spectra of the residual images, which are derived from subtracting the degraded images from the ground truth. Right, the diagram presents the mean values of the Fourier spectra based on the square length indicated on the x-axis. It is evident that varying degradations focus on distinct frequency subbands.
@@ 2.1 Hyperspectral Image Restoration
However, due to multiple or composite degradations that jointly affect structural information and fine-grained details of a given HSI, restoration methods are easily compromised, consistently leading to suboptimal performance.
@@ 2.2 All-in-One RGB Image Restoration
A recently emerging paradigm in RGB image restoration involves designing single models to restore different types of degradation [12, 7, 23, 39, 10, 5, 13]. Current techniques, as a type of multi-task learning, utilize various frameworks to manage inputs and outputs for multiple tasks with prompting methods, facilitating effective information sharing between them and corresponding degradation. Specifically, AirNet [23], the first proposed AiO image restoration method, employed contrastive learning to train the encoder to learn degradation representations and restoration with a single network. PromptIR [39] introduced the visual prompt for the degradation perception and information integration within the decoding process without a priori needs. InstructIR [10] followed the text-prompt as instruction to guide the network, archiving great results with controllability. HAIR [5] is based on the hypernetwork, which aims to generate network parameters for corresponding degradations. Recently, OneRestore [13] has extended AiO restoration from multiple degradation to composite degradation, which can address the various corruption even they exists in a same image simultaneously.
Although these methods are effective, these solutions are limited in HSIs. Due to domain gap and composite degradation, it is difficult to effectively learn the compact representation of degradation and perform effective restoration.
@@ 3 Preliminary and Motivation
3.1 Unified HSI Restoration Network Design
For comprehensive restoration using a single network, employing an architecture suitable for various degradations is crucial [8, 47]. In RGB image restoration, two classic architectures predominate: ❶ U-shape encoder-decoder [46, 45] and ❷ plain residual-in-residual [27, 17, 50]. The former excels at capturing general representations and information, being widely used in denoising, deblurring, and inpainting, but pays less attention to fine details or repeat patterns. The latter specializes in capturing detailed spatial features at the fine-grained level, serving as the backbone for super-resolution tasks [9, 41, 51].
Previous methods have overlooked the compatibility between these two restoration models, and current AiO restoration approaches [39, 5, 13] typically focus on the former rather than integrate the joint strength of both types of network design, limiting their versatility. Besides, this limitation also extends to HSI restoration models [33, 19, 21]. Therefore, we integrate a plain residual-in-residual block into the U-net cascade decoder in the PromptHSI to maintain the advantages of both designs without compromising. The versatility of the network can be ensured accordingly.
@@ 3.2 Degradations Analysis in Frequency Domain
Unlike RGB image restoration, HSI restoration requires careful consideration of spatial-spectral correlations across hundreds of bands. Our key insight comes from analyzing how different types of degradation manifest in the frequency-domain through Fourier analysis. Given a clean HSI $I_{\text{clean}}$, its Fourier transform can be expressed as:
$F(I_{\text{clean}})=F_{s}(I_{\text{clean}})+iF_{c}(I_{\text{clean}})$ (1)
where $F_{s}$ and $F_{c}$ represent the sine and cosine components respectively. Through empirical analysis, as shown in Figure 3, we observed that different types of degradation exhibit distinct patterns in how they affect these frequency components.
Specifically, the relationship between degraded HSI $I_{\text{degraded}}$ and clean HSI $I_{\text{clean}}$ in frequency-domain can be characterized by an affine transformation:
$F(I_{\text{degraded}})=\hat{\lambda}\odot F(I_{\text{clean}})+\hat{\mu}=(1+\lambda)\odot F(I_{\text{clean}})+\mu$ (2)
where $\odot$ denotes element-wise multiplication, $\lambda$ and $\mu$ are degradation-specific parameters that: $\lambda$ represents the scaling factor that modulates frequency amplitudes, $\mu$ is the bias term that shifts the frequency distribution.
Different types of degradation show characteristic patterns in $\lambda$ and $\mu$, such as ❶ Blur mainly affects high frequencies: $\lambda_{high}$ $\ll$ $\mu_{low}$. ❷ Noise adds uniform perturbations: $\mu_{noise}$approx constant $c$. ❸ Missing information creates structured patterns in both $\lambda$ and $\mu$.
@@ 4 Proposed Method
Figure 4: The architecture of the proposed PromptHSI, which is designed to tackle the composite degradations within a single universal model. First, we design a network based on U-net-like encoder-decoder architecture [46, 45] for capturing global spatial-spectral information, while combining Residual-Dense Group (RDG) and Fast-Residual-Dense Block (FRDB) presented in DRCT [17] and RTCS [16], respectively, in the proposed PGFAB for spatial-spectral enhancement. Afterwards, text-prompt-guidance and frequency-aware feature map modulation is proposed to jointly better capture the representation of composite degradation with controllability.
@@ 4.1 Hyperspectral Image Restoration Pipeline
Spatial and spectral feature enhancement. The output of PGFM is first improved by consecutive RDG [17] and FRDB [16] to improve fine-grained details. Then these features are interacted by cross-attention layer, which aims to comprehensively integrate the spatial-spectral feature information from different branches across all dimensions. Specifically, it can be expressed as follows:
$\{Q_{\alpha},K_{\alpha},V_{\alpha}\}=\mathcal{F}^{\alpha}_{qkv}(F_{\alpha}),\quad\{Q_{\beta},K_{\beta},V_{\beta}\}=\mathcal{F}^{\beta}_{qkv}(F_{\beta}),$ (3)
where $F_{\alpha},F_{\beta}$ denote features from the spatial and spectral enhancement block. Subsequently, we exchange the queries $Q$ of two branches for spatial-spectral interaction:
$F_{f}^{\beta}=softmax(\frac{Q_{\alpha}K_{\beta}}{d_{k}})V_{\beta},\quad F_{f}^{\alpha}=softmax(\frac{Q_{\beta}K_{\alpha}}{d_{k}})V_{\alpha},$ (4)
where $d_{k}$ is the scaling factor. Subsequently, we concatenate the results obtained by the cross-attention through $F_{f}^{\gamma}=Concat({F_{f}^{\beta},F_{f}^{\alpha}})$ to obtain the integrated features.
$\{Q_{\gamma},K_{\gamma},V_{\gamma}\}=\mathcal{F}^{\gamma}_{qkv}(F_{f}^{\gamma}),\quad\hat{F_{f}^{\gamma}}=softmax(\frac{Q_{\gamma}K_{\gamma}}{d_{k}})V_{\gamma},$ (5)
Finally, we employ spectral-spatial self-attention for feature co-enhancement to improve representation capability.
Through a dual-branch architecture, we integrate text-prompt guidance effectively across both spatial and spectral dimensions. By incorporating plain-residual-in-residual [27, 17, 16] modules within the U-shape cascade decoder [46, 8], we enhance fine-grained image restoration while integrating global information, thereby effectively addressing various degradations for HSI restoration.

### v3

# PromptHSI: Universal Hyperspectral Image Restoration with Vision-Language Modulated Frequency Adaptation

@@ Abstract
Recent advances in All-in-One (AiO) RGB image restoration have demonstrated the effectiveness of prompt learning in handling multiple degradations within a single model. However, extending these approaches to hyperspectral image (HSI) restoration is challenging due to the domain gap between RGB and HSI features, information loss in visual prompts under severe composite degradations, and difficulties in capturing HSI-specific degradation patterns via text prompts. In this paper, we propose PromptHSI, the first universal AiO HSI restoration framework that addresses these challenges. By incorporating frequency-aware feature modulation—which utilizes frequency analysis to narrow down the restoration search space—and employing vision-language model (VLM)-guided prompt learning, our approach decomposes text prompts into intensity and bias controllers that effectively guide the restoration process while mitigating domain discrepancies. Extensive experiments demonstrate that our unified architecture excels at both fine-grained recovery and global information restoration across diverse degradation scenarios, highlighting its significant potential for practical remote sensing applications. The source code is available at https://github.com/chingheng0808/PromptHSI.
Corresponding author, email: cchsu@gs.ncku.edu.twProject page: https://chingheng0808.github.io/prompthsiP/static.html
@@ 1 Introduction
Figure 1: Performance comparison between All-in-One and HSI restoration methods on the AVIRIS dataset. The proposed PromptHSI offers leading performance and controllability while retaining a compact model size.
@@
Hyperspectral images encompass a greater number of spectral bands than RGB images, offering extensive spectral details [52]. Because materials reflect different spectral bands, each substance can have a unique signature for material identification [1, 45, 52, 38]. HSI can capture this valuable information and has broad application value and potential. However, because of the limitation of the device and bandwidth, observed HSIs are usually damaged by different types of degradation from camera sensing and transmission. Thus, HSI restoration is a crucial pre-processing step to recover degraded pixels from flawed observations.
To address these issues, low-rank tensor models [49, 42] are widely used in hyperspectral image restoration, relying on the assumption that the clean image $\mathbf{X}$ exhibits low-rank properties [50, 13] across spatial and spectral dimensions. The restoration process is often formulated as $\mathbf{X}\approx\mathbf{L}+\mathbf{S}+\mathbf{N}$, where $\mathbf{L}$ represents the underlying low-rank structure of the clean image, $\mathbf{S}$ denotes structured sparse noise (e.g., stripe noise or dead pixels), $\mathbf{N}$ represents random noise, such as Gaussian or Poisson noise. However, under multiple degradations, this assumption often fails. Structured noise components $\mathbf{S}$ introduce high-rank artifacts, increasing the observed rank: $\operatorname{rank}(\mathbf{X})\gg\operatorname{rank}(\mathbf{L})$, which makes low-rank approximations less accurate. Furthermore, mixed noise complicates the separation between the clean component $\mathbf{L}$ and noise components $\mathbf{S}$ and $\mathbf{N}$, often leading to over-smoothing:
$\min_{\mathbf{L}}\|\mathcal{P}_{\Omega}(\mathbf{X}-\mathbf{L})\|_{F}^{2}+\lambda\operatorname{TV}(\mathbf{L}),$ (1)
where $P_{\Omega}$ is a projection operator on observed pixels, and $\operatorname{TV}(\mathbf{L})$ is a total variation regularization term. While this enforces spatial smoothness, it often suppresses fine details, degrading spectral fidelity. Therefore, when multiple degradations are present, the core low-rank assumption is violated, making traditional tensor-based approaches ineffective. More adaptive methods that incorporate hybrid priors [34, 33, 63] or deep learning-based corrections are needed to achieve robust restoration.
The emergence of AiO RGB image restoration methods [47, 4, 36, 10, 15, 11] has demonstrated remarkable success in handling diverse degradations within a unified network. These methods leverage prompt learning, either through visual prompts [47, 4, 11] or text-based prompts [10, 15], to adaptively restore corrupted images. However, despite their effectiveness in RGB restoration, these AiO approaches face fundamental limitations when extended to HSI restoration. The primary challenge arises from the inherent complexity and variability of degradation patterns in HSI data, making it difficult to generalize the same prompt-based strategies.
To address these challenges, we propose PromptHSI, a frequency-aware, prompt-guided HSI restoration framework that overcomes the limitations of existing AiO RGB-based restoration methods. Unlike conventional approaches that struggle with the complexity of HSI degradation patterns, PromptHSI integrates frequency-domain priors with VLM-driven prompt learning, enabling robust and adaptive restoration across diverse degradation scenarios. In summary, our key contributions are summarized as follows:
  • A Frequency-Aware Prior for Efficient Restoration. We introduce a structured frequency-domain analysis that models the impact of different degradations on specific frequency components. By leveraging Fourier-based priors, PromptHSI significantly reduces the restoration search space, making it easier for deep networks to learn effective recovery mappings.
  • Vision-Language Modulated Prompt Learning for Controllable Restoration. Unlike traditional prompt-based approaches that rely on static feature modulation, PromptHSI employs VLM-guided adaptive modulation, where textual embeddings dynamically control the restoration process. This allows for more flexible and interpretable restoration outcomes, surpassing conventional AiO RGB methods in handling complex degradations.
  • The First Composite Degradation Dataset for Benchmarking HSI Restoration. To facilitate research in HSI restoration, we construct the first standardized composite degradation dataset, incorporating cloud occlusion, blurring, noise, and spectral band loss. We believe this dataset will be publicly released to support future developments in hyperspectral image restoration.
Through extensive experiments, we demonstrate that PromptHSI consistently outperforms state-of-the-art AiO RGB methods in HSI restoration tasks while offering controllable restoration via text prompts.
@@
Figure 2: Illustration of the degraded HSI and description synthesis pipeline. Each degradation is activated with probability $p$ (set to 0.5 in all experiments). To ensure diversity and realism, each activated degradation randomly applies one of its sub-degradations to the input HSI. More details are provided in the Supplementary.
@@ 2.1 Hyperspectral Image Restoration
Despite these advancements, multiple or composite degradations simultaneously affect both structural integrity and fine-grained details in HSIs, making restoration methods highly prone to performance degradation. Moreover, erroneous compensation effects [37, 27] further exacerbate this issue, as models may misinterpret different degradation types, leading to incorrect adjustments. For instance, in HSIs degraded by both noise and cloud occlusion, conventional denoising methods may mistakenly treat occluded regions as random noise, resulting in spectral distortions and suboptimal restoration outcomes.
@@ 2.2 All-in-One RGB Image Restoration
A recently emerging paradigm in RGB image restoration designs single models to handle multiple degradations [14, 6, 27, 47, 10, 4, 15]. These multi-task methods leverage prompting and diverse frameworks to share information between tasks and degradations. For example, AirNet [27] used contrastive learning to jointly learn degradation representations and restoration; PromptIR [47] introduced visual prompts for degradation perception during decoding; InstructIR [10] employed text prompts to guide the network; HAIR [4] utilized a hypernetwork to generate degradation-specific parameters; OneRestore [15] extended AiO restoration to composite degradations, addressing multiple corruptions in a single image; and AdaIR [11] transforms feature maps into the frequency domain via Fourier transform and uses an adaptive mask, guided by the degraded image, to decouple low- and high-frequency features.
However, despite their effectiveness in RGB, these solutions are limited in HSIs. Due to domain gaps and composite degradation, learning a compact degradation representation and achieving effective restoration in HSIs remains challenging.
@@ 3 Preliminary and Motivation
Figure 3: Visualizations of the Fourier transforms of HSIs with specific degradations, displayed from top to bottom alongside the Fourier spectra of the residual images (obtained by subtracting the degraded images from the ground truth).
3.1 Hybrid Architecture for HSI Restoration
Effective HSI restoration requires an architecture capable of addressing diverse degradations within a single framework [8, 55]. Existing approaches predominantly adopt (i) U-shaped encoder-decoder networks [54, 53], which effectively capture global semantic representations for overall restoration but have difficulty with fine details, or (ii) plain residual-in-residual networks [31, 19, 62], which enhance spatial features and high-frequency information but lack global awareness.
AiO restoration models [47, 4, 15] and HSI-specific methods [39, 21, 24] favor the former, limiting fine-detail recovery. To overcome this, PromptHSI integrates residual-in-residual blocks within a U-shaped cascade decoder, preserving global context while refining local details for greater adaptability across degradation scenarios.
@@ 3.2 Degradations Analysis in Frequency Domain
Unlike RGB image restoration, HSI restoration must account for complex spatial-spectral correlations across hundreds of spectral bands [58]. The degradation process can be mathematically formulated as $I_{\text{degraded}}=\mathcal{D}(I_{\text{clean}})+\epsilon$, where $\mathcal{D}(\cdot)$ represents the degradation operator, encompassing effects such as blurring, stripe noise, and occlusion, while $\epsilon$ denotes additive noise.
Frequency modulation. Recovering $I_{\text{degraded}}$ from an arbitrary degradation function $\mathcal{D}^{-1}\{\cdot\}$ is inherently challenging, particularly in the presence of composite degradations, where multiple degradation factors (e.g., noise, blur, and missing bands) interact in a complex manner. Unlike isolated degradations, which can often be addressed with predefined priors, composite degradations induce coupled distortions across both spatial and spectral dimensions, making traditional inverse modeling impractical.
Our empirical analysis (refer to Figure 3) reveals that different types of degradation exhibit distinct but overlapping spectral characteristics. Motivated by this, we reformulate Equation 2 by modeling the degradation process as an affine transformation in the frequency domain:
$\footnotesize F(I_{\text{degraded}})=\hat{\lambda}\odot F(I_{\text{clean}})+\hat{\mu}=(1+\lambda)\odot F(I_{\text{clean}})+\mu,$ (2)
where $F(\cdot)$ denotes the Fourier transform, $\lambda$ modulates frequency amplitudes, $\mu$ shifts the frequency distribution, and $\odot$ represents element-wise multiplication. Under composite degradations, these frequency parameters exhibit complex dependencies:
  • Blurring attenuates high-frequency components ($\lambda_{\text{blur}}\ll\mu_{\text{blur}}$), reducing fine details.
  • Noise induces broadband perturbations, where ($\mu_{\text{noise}}\approx constant$) remains nearly uniform across frequencies.
  • Occlusion and missing bands cause structured disruptions in both $\lambda$ and $\mu$, leading to spectral inconsistencies.
  • Composite degradations result in frequency interactions, where overlapping effects create non-trivial distortions beyond individual degradations.
For a clearer understanding, Figure 4 (left) visualizes these degradation-specific frequency characteristics, highlighting the non-additive nature of composite degradations in the Fourier domain.
Figure 4: A diagram illustrating our proposed strategy, where integrating prompt guidance with frequency modulation facilitates the adaptive selection of $\lambda$ and $\mu$, enhancing HSI restoration under various degradation types. This mechanism encourages the model to efficiently converge toward optimal parameters.
Figure 5: The architecture of the proposed PromptHSI, which is designed to tackle the composite degradations within a single universal model via VLM-guided frequency modulation. First, we design a network based on U-net-like encoder-decoder architecture for capturing global spatial-spectral information, while combining plain-residual-in residual networks in the proposed PGFAB for spatial-spectral enhancement. Afterwards, text-prompt-guidance and frequency-aware feature modulation is proposed to jointly better capture the representation of composite degradation with controllability.
Prompt guidance. The diverse nature of HSI degradations presents significant challenges in learning the corresponding frequency coefficients. As shown in Figure 4 (right), previous AiO restoration methods [4, 11] utilize adaptive modules to classify degradation types. However, lacking explicit prompt-based guidance, these methods rely solely on implicit feature learning, making parameter selection unreliable, particularly under composite degradations, where multiple degradation effects exhibit complex spectral interactions.
To address this, we introduce a VLM-driven prompt guidance mechanism, a core component of PromptHSI, which embeds frequency-aware priors into restoration. Leveraging textual cues from a given prompt $T$, our approach modulates frequency coefficients, refining the parameter space for faster convergence and enhanced frequency adaptation. This ensures precise control, adaptability, and robust performance across diverse degradations.
@@ 4.1 Hyperspectral Image Restoration Pipeline
Spatial and spectral feature enhancement. The PGFM output is first refined using sequential RDG [19] and FRDB [18] modules to enhance fine details. Next, a cross-attention layer integrates spatial and spectral features from different branches. Specifically, we compute:
$\footnotesize\{Q_{\alpha},K_{\alpha},V_{\alpha}\}=\mathcal{F}^{\alpha}_{qkv}(F_{\alpha}),\quad\{Q_{\beta},K_{\beta},V_{\beta}\}=\mathcal{F}^{\beta}_{qkv}(F_{\beta}),$ (3)
where $F_{\alpha}$ and $F_{\beta}$ are the spatial and spectral features, respectively. We then exchange queries between branches:
$\footnotesize F_{f}^{\beta}=\operatorname{softmax}\left(\frac{Q_{\alpha}K_{\beta}}{d_{k}}\right)V_{\beta},\quad F_{f}^{\alpha}=\operatorname{softmax}\left(\frac{Q_{\beta}K_{\alpha}}{d_{k}}\right)V_{\alpha},$ (4)
where $d_{k}$ is a scaling factor. The outputs $F_{f}^{\alpha}$ and $F_{f}^{\beta}$ are combined into $F_{f}^{\gamma}$, which is then enhanced through cross-attention for feature intercalation:
$\footnotesize\{Q_{\gamma},K_{\gamma},V_{\gamma}\}=\mathcal{F}^{\gamma}_{qkv}(F_{f}^{\gamma}),\quad\hat{F}_{f}^{\gamma}=\operatorname{softmax}\left(\frac{Q_{\gamma}K_{\gamma}}{d_{k}}\right)V_{\gamma}.$ (5)
Our dual-branch architecture effectively integrates text-prompt guidance across spatial and spectral dimensions. By embedding plain-residual-in-residual modules [31, 19, 18] within a U-shaped cascade decoder [54, 8], our framework achieves both fine-grained and global restoration, effectively addressing various HSI degradations.